OpenStudy (anonymous):

some one please help me!!! WILL GIVE MEDAL FOR HELP! f(x)=2-4x what is f(-2) A) -6 B) 1 C) 4 D) 10

OpenStudy (anonymous):

\[f(-2)=2-4(-2)=2+8=10\]

OpenStudy (akashdeepdeb):

If f(x) = 2-4x Then f(a) = 2-4a f(b) = 2-4b So f(-2) = 2 - 4(-2) = 2+8 Understood? :)

OpenStudy (anonymous):

thank yall!!!